# Asymptotic evaluation of the information processing capacity in reservoir computing

### v1

@@ 3 Asymptotic expansion of the IPC
[…]
Following Ref. [14], the asymptotic forms can be derived from the mean-value theorem. From the mean-value theorem, there exists $c\in[0,1]$ which satisfies
$0=$ $\nabla_{w}l_{T}(w_{T})=\nabla_{w}l_{T}(w_{0})+\nabla_{w}^{2}l_{T}(w^{\prime})\,(w_{T}-w_{0}),$ (28)
$w^{\prime}=$ $cw_{T}+(1-c)w_{0}.$ (29)
Substituting $\nabla_{w}l_{T}(w_{0})=2(J/T)^{1/2}\xi_{T}$ and $\nabla_{w}^{2}l_{T}(w^{\prime})=J+o_{p}(1)$ into Eq. (28), we obtain the asymptotic form of $w_{T}$,
$w_{T}=w_{0}-(TJ)^{-1/2}\,\xi_{T}+o_{p}\left(\frac{1}{\sqrt{T}}\right).$ (30)
The third and higher order derivatives of $l_{T}$ is 00. Hence, applying the mean-value theorem to Eqs. (11), we find that
$l_{T}(w_{0})=$ $l_{T}(w_{T})+\nabla_{w}l_{T}(w_{T})^{\top}(w_{0}-w_{T})+\frac{1}{2}(w_{0}-w_{T})^{\top}\,\nabla_{w}^{2}l_{T}(w_{T})\,(w_{0}-w_{T}).$ (31)
Substituting $\nabla_{w}l_{T}(w_{T})=0$, $\nabla_{w}^{2}l_{T}(w_{T})=2J$ and Eq. (30) into Eq. (31), we obtain
$l_{T}(w_{T})=l_{T}(w_{0})-\frac{1}{T}\|\xi_{T}\|^{2}+o_{p}\left(\frac{1}{T}\right).$ (32)
Similarly, applying the mean-value theorem to Eq. (12), we find that
$l^{\prime}_{T^{\prime}}(w_{T})=$ $l^{\prime}_{T^{\prime}}(w_{0})+\nabla_{w}l^{\prime}_{T^{\prime}}(w_{0})^{\top}(w_{T}-w_{0})+\frac{1}{2}(w_{T}-w_{0})^{\top}\,\nabla_{w}^{2}l^{\prime}_{T^{\prime}}(w_{0})\,(w_{T}-w_{0}).$ (33)
Substituting $\nabla_{w}l^{\prime}_{T^{\prime}}(w_{0})=2(J/T^{\prime})^{1/2}\xi^{\prime}_{T^{\prime}}+o_{p}(1/\sqrt{T^{\prime}})$, $\nabla_{w}^{2}l^{\prime}_{T^{\prime}}(w_{0})=2J$ and Eq. (30) into Eq. (33), we obtain
$l^{\prime}_{T^{\prime}}(w_{T})=$ $l^{\prime}_{T^{\prime}}(w_{0})+\frac{1}{T}\|\xi_{T}\|^{2}+\frac{2}{\sqrt{TT^{\prime}}}\,\xi_{T}^{\top}\,\xi^{\prime}_{T^{\prime}}+o_{p}\left(\frac{1}{T}\right)\,.$ (34)
From the central limit theorem [15], the denominators of Eqs. (9) and (10) converge in distribution,
[…]
Substituting Eqs. (32), (34) and (35) into Eqs. (9) and (10), we obtain the asymptotic forms of the training and the test IPC,
$C_{T}=$ $1-\frac{1}{\mu_{0}}\left[l_{T}(w_{0})-\frac{l_{T}(w_{0})\zeta_{T}}{\mu_{0}}+\frac{l_{T}(w_{0})\zeta_{T}^{2}}{\mu_{0}^{2}}-\frac{\|\xi_{T}\|^{2}}{T}\right]+o_{p}\left(\frac{1}{T}\right),$ (36)
$C^{\prime}_{T,T^{\prime}}=$ $1-\frac{1}{\mu_{0}}\left[l^{\prime}_{T^{\prime}}(w_{0})-\frac{l^{\prime}_{T^{\prime}}(w_{0})\zeta_{T^{\prime}}}{\mu_{0}}+\frac{l^{\prime}_{T^{\prime}}(w_{0})\zeta_{T^{\prime}}^{2}}{\mu_{0}^{2}}+\frac{\|\xi_{T}\|^{2}}{T}+\frac{2\xi_{T}^{\top}\xi^{\prime}_{T^{\prime}}}{\sqrt{TT^{\prime}}}\right]+o_{p}\left(\frac{1}{T}\right).$ (37)
@@
Taking expectations for the training and the test data, the mean and the variance of the training IPC are given by
$\mathbb{E}[C_{T}]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{C_{l,\zeta}}{\mu_{0}^{2}}-\frac{l(w_{0})\,V_{\zeta,\infty}}{\ mu_{0}^{3}}+\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,,$ (38)
$\mathbb{V}[C_{T}]=$ $\frac{1}{T}\left[\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{l(w_{0})^{2}\,V_{\zeta,\infty}}{\mu_{0}^{4}}-\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\zeta}\right]+o\left(\frac{1}{T}\right)\,.$ (39)
Similarly, the mean and the variance of the test IPC are given by
$\mathbb{E}[C^{\prime}_{T,T^{\prime}}]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{C_{l,\zeta}}{\mu_{0}^{2}}-\frac{T}{T^{\prime}}\,\frac{l(w_{0})\,V_{\zeta,\infty}}{\mu_{0}^{3}}-\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,,$ (40)
$\mathbb{V}[C^{\prime}_{T,T^{\prime}}]=$ $\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{T}{T^{\prime}}\,\frac{l(w_{0})^{2}\,V_{\zeta,\infty}}{\mu_{0}^{4}}-\frac{T}{T^{\prime}}\,\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\zeta}\right]+o\left(\frac{1}{T}\right)\,.$ (41)
@@
Note that we have assumed $T^{\prime}=O(T)$. From Eqs. (38) and (40), we find that both means approach the true IPC, $C_{0}=1-l(w_{0})/\mu_{0}$, in $O(1/T)$. To estimate the true IPC, we first approximate the expectation values, $\mathbb{E}[C_{T}]$ and $\mathbb{E}[C^{\prime}_{T,T^{\prime}}]$, by the sample mean at each $T$. Then we ignore the term $o(1/T)$ in Eqs. (38) and (40) and use the least squares method to estimate the true IPC (see Appendix). Our method can be applied to RC systems that satisfy the assumption that $U$ is a stationary ergodic process, $\mathbb{V}[l(w,U^{\prime})]<\infty$, and $I_{\infty}(w),J,V_{\zeta,\infty},C_{l,\zeta}(w)$, and $V_{l,\infty}(w)$ are finite matrices. In the next section, we apply this method to several models.
@@
Finally, we note that fitting the variances by Eqs. (39) and (41) fails when the true IPC is 00. In this case, the true parameter is $w_{0}=0$, and we find $l(w_{0},U^{\prime})=\|\hat{y}(U^{\prime})\|^{2}$. Thus, we obtain
${\rm Cov}(l(w_{0},U^{\prime}_{0}),l(w_{0},U^{\prime}_{t}))={\rm Cov}(l(w_{0},U^{\prime}_{0}),\|\hat{y}(U^{\prime}_{t})\|^{2})={\rm Cov}(l(w_{0},U^{\prime}_{0}),\zeta(U^{\prime}_{t}))={\rm Cov}(\zeta(U^{\prime}_{0}),\zeta(U^{\prime}_{t})).$ (42)
That is, $V_{l,\infty}=V_{\zeta,\infty}=C_{l,\zeta}$. Since the coefficients of the $1/T$ terms in Eqs. (39) and (41) vanish, fitting the coefficients of the $1/T$ terms fails. In the next section, we will show an example of the failure.

### v2

@@ 3.1 Asymptotic expansion of the IPC
$\xi_{T}\xrightarrow{d}$ ${\cal N}(0,J^{-1/2}I_{\infty}J^{-1/2}),$ (30)
$\xi^{\prime}_{T^{\prime}}\xrightarrow{d}$ ${\cal N}(0,J^{-1/2}I_{\infty}J^{-1/2}).$ (31)
Following Ref. [watanabe2009algebraic], we expand $\nabla_{w}\bar{l}_{T}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})$ around $w_{0}$ and obtain
$0=$ $\nabla_{w}\bar{l}_{T}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})=\nabla_{w}\bar{l}_{T}(w_{0},U^{\prime}_{T})+\nabla_{w}^{2}\bar{l}_{T}(U^{\prime}_{T})\,(w_{T}(U^{\prime}_{T})-w_{0}),$ (32)
$\nabla_{w}\bar{l}_{T}(w_{0},U^{\prime}_{T})=$ $-\frac{2}{T}\sum_{t=1}^{T}(\hat{y}(U^{\prime}_{t})-w_{0}^{\top}x(U^{\prime}_{t}))\,x(U^{\prime}_{t})^{\top},$ (33)
$\nabla_{w}^{2}\bar{l}_{T}(U^{\prime}_{T})=$ $\frac{2}{T}\sum_{t=1}^{T}x(U^{\prime}_{t})\,x(U^{\prime}_{t})^{\top}.$ (34)
Substituting $\nabla_{w}l_{T}(w_{0},U^{\prime}_{T})=2(J/T)^{1/2}\xi_{T}$ and $\nabla_{w}^{2}\bar{l}_{T}(U^{\prime}_{T})=2J+o_{p}(1)$ into Eq. (32), we find the asymptotic form of $w_{T}$,
$w_{T}(U^{\prime}_{T})=w_{0}-(TJ)^{-1/2}\,\xi_{T}+o_{p}\left(\frac{1}{\sqrt{T}}\right).$ (35)
Hence, expanding Eq. (12) around $w_{T}(U^{\prime}_{T})$, we find that
$\bar{l}_{T}(w_{0},U^{\prime}_{T})=$ $\bar{l}_{T}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})+\nabla_{w}{\bar{l}_{T}}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})^{\top}(w_{0}-w_{T})$
$+\frac{1}{2}(w_{0}-w_{T}(U^{\prime}_{T}))^{\top}\,\nabla_{w}^{2}\bar{l}_{T}(U^{\prime}_{T})\,(w_{0}-w_{T}(U^{\prime}_{T})).$ (36)
Substituting $\nabla_{w}\bar{l}_{T}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})=0$, $\nabla_{w}^{2}\bar{l}_{T}(U^{\prime}_{T})=2J+o_{p}(1)$ and Eq. (35) into Eq. (36), we obtain
$\bar{l}_{T}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T})=\bar{l}_{T}(w_{0},U^{\prime}_{T})-\frac{1}{T}\xi_{T}^{2}+o_{p}\left(\frac{1}{T}\right).$ (37)
Similarly, expanding Eq. (13) around $w_{0}$, we find that
$\bar{l}^{\prime}_{T,T^{\prime}}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T+T^{\prime}})=$ $\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})+\nabla_{w}\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})^{\top}(w_{T}(U^{\prime}_{T})-w_{0})$
$+\frac{1}{2}(w_{T}(U^{\prime}_{T})-w_{0})^{\top}\,\nabla_{w}^{2}\bar{l}^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})\,(w_{T}(U^{\prime}_{T})-w_{0}).$ (38)
Substituting $\nabla_{w}\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})=2(J/T^{\prime})^{1/2}\xi^{\prime}_{T^{\prime}}$, $\nabla_{w}^{2}\bar{l}^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})=2J+o_{p}(1)$ and Eq. (35) into Eq. (38), we obtain
$\bar{l}^{\prime}_{T,T^{\prime}}(w_{T}(U^{\prime}_{T}),U^{\prime}_{T+T^{\prime}})=$ $\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})+\frac{1}{T}(\xi_{T})^{2}+\frac{2}{\sqrt{TT^{\prime}}}\,\xi_{T}^{\top}\,\xi^{\prime}_{T^{\prime}}+o_{p}\left(\frac{1}{T}\right)\,.$ (39)
Now, substituting Eqs. (37), (39) and (24) into Eqs. (10) and (11), we obtain the asymptotic forms of the training and the test IPC,
$C_{T}(U^{\prime}_{T})$
$=$ $\left(1-\frac{\bar{l}_{T}(w_{0},U^{\prime}_{T})}{\bar{\mu}_{T}(U^{\prime}_{T})}\right)+\frac{1}{T}\frac{(\xi_{T})^{2}}{\bar{\mu}_{T}(U^{\prime}_{T})}+o_{p}\left(\frac{1}{T}\right)$ (40)
$=$ $1-\frac{1}{\mu_{0}}\left[\bar{l}_{T}(w_{0},U^{\prime}_{T})-\frac{\bar{l}_{T}(w_{0},U^{\prime}_{T})\zeta_{T}}{\mu_{0}}+\frac{\bar{l}_{T}(w_{0},U^{\prime}_{T})\zeta_{T}^{2}}{\mu_{0}^{2}}-\frac{(\xi_{T})^{2}}{T}\right]+o_{p}\left(\frac{1}{T}\right),$ (41)
$C^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})$
$=$ $\left(1-\frac{\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})}{\bar{\mu}_{T^{\prime}}(U^{\prime}_{T+T^{\prime}})}\right)-\frac{1}{T}\left(\frac{(\xi_{T})^{2}}{\bar{\mu}_{T^{\prime}}(U_{T+T^{\prime}})}+\sqrt{\frac{T}{T^{\prime}}}\frac{2\xi_{T}^{\top}\xi^{\prime}_{T^{\prime}}}{\bar{\mu}_{T^{\prime}}(U^{\prime}_{T+T^{\prime}})}\right)+o_{p}\left(\frac{1}{T}\right)$ (42)
$=$ $1-\frac{1}{\mu_{0}}\Biggl[\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})-\frac{\bar{l}^{\prime}_{T,T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})\zeta_{T^{\prime}}}{\mu_{0}}+\frac{\bar{l}^{\prime}_{T^{\prime}}(w_{0},U^{\prime}_{T+T^{\prime}})\zeta_{T^{\prime}}^{2}}{\mu_{0}^{2}}$
$+\frac{(\xi_{T})^{2}}{T}+\frac{2\xi_{T}^{\top}\xi^{\prime}_{T^{\prime}}}{\sqrt{TT^{\prime}}}\Biggr]+o_{p}\left(\frac{1}{T}\right),$ (43)
$\zeta_{T}\equiv$ $\bar{\mu}_{T}(U^{\prime}_{T})-\mu_{0},\ \ \zeta_{T^{\prime}}\equiv\bar{\mu}_{T^{\prime}}(U^{\prime}_{T+T^{\prime}})-\mu_{0}.$ (44)
Note that we have assumed $T^{\prime}=O(T)$. The first and the rest terms of Eq. (40) represent the model assumption bias and the finite-size effect, respectively. If the RC system can reproduce the target function $\hat{y}$, the first term converges to $1$ in $T\to\infty$. The first and the rest terms of Eq. (42) can similarly be interpreted.
@@
Taking expectations for the training and the test data in Eqs. (41), the mean and the variance of the training IPC are given by
$\mathbb{E}[C_{T}(U^{\prime}_{T})]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{C_{l,\mu}}{\mu_{0}^{2}}-\frac{l(w_{0})\,V_{\mu,\infty}}{\mu_{0}^{3}}+\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,$, (45)
$\mathbb{V}[C_{T}(U^{\prime}_{T})]=$ $\frac{1}{T}\left[\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{l(w_{0})^{2}\,V_{\mu,\infty}}{\mu_{0}^{4}}-\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\mu}\right]+o\left(\frac{1}{T}\right)\,.$ (46)
Similarly, from Eq. (43), the mean and variance of the test IPC are given by
$\mathbb{E}[C^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})]=$ $1-\frac{l(w_{0})}{\mu_{0}}+\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{C_{l,\mu}}{\mu_{0}^{2}}-\frac{T}{T^{\prime}}\,\frac{l(w_{0})\,V_{\mu,\infty}}{\mu_{0}^{3}}-\frac{{\rm Tr}(I_{\infty}J^{-1})}{\mu_{0}}\right]+o\left(\frac{1}{T}\right)\,$, (47)
$\mathbb{V}[C^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})]=$ $\frac{1}{T}\left[\frac{T}{T^{\prime}}\,\frac{V_{l,\infty}}{\mu_{0}^{2}}+\frac{T}{T^{\prime}}\,\frac{l(w_{0})^{2}\,V_{\mu,\infty}}{\mu_{0}^{4}}-\frac{T}{T^{\prime}}\,\frac{2}{\mu_{0}^{3}}\,l(w_{0})\,C_{l,\mu}\right]+o\left(\frac{1}{T}\right)\,.$ (48)
@@
From Eqs. (45) and (47), we find that both means approach the true IPC, $C_{0}=1-l(w_{0})/\mu_{0}$, in $O(1/T)$. To estimate the true IPC, we first approximate the expectation values, $\mathbb{E}[C_{T}(U^{\prime}_{T})]$ and $\mathbb{E}[C^{\prime}_{T,T^{\prime}}(U^{\prime}_{T+T^{\prime}})]$, by the sample mean at each $T$. Then we ignore the term $o(1/T)$ in Eqs. (45) and (47) and use the weighted least squares method to estimate the true IPC (see Appendix).
@@
Note that fitting the variances by Eqs. (46) and (48) fails when the true IPC is 0. In this case, the true parameter is $w_{0}=0$, and we find $l(w_{0},U^{\prime})=(\hat{y}(U^{\prime}))^{2}$. Thus, we obtain
${\rm Cov}(l(w_{0},U^{\prime}_{0}),l(w_{0},U^{\prime}_{t}))={\rm Cov}(l(w_{0},U^{\prime}_{0}),(\hat{y}(U^{\prime}_{t}))^{2})$
$=$ ${\rm Cov}(l(w_{0},U^{\prime}_{0}),\mu(U^{\prime}_{t}))={\rm Cov}(\mu(U^{\prime}_{0}),\mu(U^{\prime}_{t})).$ (49)
That is, $V_{l,\infty}=V_{\mu,\infty}=C_{l,\mu}$. Since the coefficients of the $1/T$ terms in Eqs. (46) and (48) vanish, fitting the coefficients of the $1/T$ terms fails. Although we do not show whether the converse holds, we use this property to determine whether the true IPC is 0. If the estimated true IPC is close to 0 and the power of the variance is not $1/T$, we propose that the true IPC is taken to be 0. In Ref. [dambre2012information], by assuming that $x(U_{t})$ and $\hat{y}(U_{t})$ are independent at each time, it is derived that when the true IPC is 0, the variance of the training IPC at $T$ is $O(1/T^{2})$, which is consistent with our result. Since discrimination by the power of the variance is less affected by fluctuations in finite-length data than discrimination by numerical value, our method is expected to be more robust.
3.2 Remarks on the proposed method
We make three comments. First, when we evaluate the training and the test IPC, we need not store $(x(u^{\prime}_{t}))_{t=1}^{T+T^{\prime}}$ and $(\hat{y}(u_{t}))_{t=1}^{T+T^{\prime}}$. Since Eqs. (9), (10), and (11) can be rewritten as
$w_{T}(u^{\prime}_{T})=\left<xx^{\top}\right>_{T}^{-1}\left<x\hat{y}^{\top}\right>_{T},$ (50)
$C_{T}(u^{\prime}_{T})=\frac{1}{\left<\hat{y}^{2}\right>_{T}}{\rm Tr}\left[\left<\hat{y}x^{\top}\right>_{T}\left<xx^{\top}\right>_{T}^{-1}\left<x\hat{y}^{\top}\right>_{T}\right],$ (51)
$C^{\prime}_{T,T^{\prime}}(u^{\prime}_{T+T^{\prime}})=\frac{1}{\left<\hat{y}^{2}\right>_{T^{\prime}}}{\rm Tr}\biggl[\left<\hat{y}x^{\top}\right>_{T}\left<xx^{\top}\right>_{T}^{-1}\left(2\left<x\hat{y}^{\top}\right>_{T^{\prime}}-\left<xx^{\top}\right>_{T^{\prime}}\left<xx^{\top}\right>_{T}^{-1}\left<x\hat{y}^{\top}\right>_{T}\right)\biggr],$ (52)
$\left<\hat{y}^{2}\right>_{T}\equiv\frac{1}{T}\sum_{t=1}^{T}\hat{y}(u^{\prime}_{t})^{2},\ \left<\hat{y}x^{\top}\right>_{T}\equiv\frac{1}{T}\sum_{t=1}^{T}\hat{y}(u^{\prime}_{t})x(u^{\prime}_{t})^{\top},\ \left<xx^{\top}\right>_{T}\equiv\frac{1}{T}\sum_{t=1}^{T}x(u^{\prime}_{t})x(u^{\prime}_{t})^{\top},$ (53)
$\left<\hat{y}^{2}\right>_{T^{\prime}}\equiv\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\hat{y}(u^{\prime}_{t})^{2},\ \left<\hat{y}x^{\top}\right>_{T^{\prime}}\equiv\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}\hat{y}(u^{\prime}_{t})x(u^{\prime}_{t})^{\top},\ \left<xx^{\top}\right>_{T^{\prime}}\equiv\frac{1}{T^{\prime}}\sum_{t=T+1}^{T+T^{\prime}}x(u^{\prime}_{t})x(u^{\prime}_{t})^{\top},$ (54)
it is only necessary to sequentially store the summand of Eqs. (53) and (54), e.g. $\sum_{t=1}^{s}\hat{y}(u^{\prime}_{t})^{2}$ at time $s$, at each time up to $T$ or $T+T^{\prime}$. The number of components of the sequences, $(x(u^{\prime}_{t}))_{t=1}^{T+T^{\prime}}$ and $(\hat{y}(u_{t}))_{t=1}^{T+T^{\prime}}$, is $(T+T^{\prime})(d_{1}+d_{2})$, and the number of components of the summand is $2(d_{1}^{2}+d_{1}d_{2}+1)$. To estimate the IPC for infinitely long data, $T+T^{\prime}>2d_{1}$ is satisfied in most cases, and therefore storing the summand instead of the sequences reduces memory consumption.
Secondly, we estimate the computational complexity of our method. The length of data should be longer than the number of parameters, $T>d_{1}$, and the number of hidden nodes is set to be larger than the input and output dimensions in general, $d_{1}\geq d_{0},d_{2}$. We assume that the training and test data are of comparable length. First, we estimate the computational complexity of the training and the test IPC for each $T$ and $T^{\prime}$. The computational complexity of $x_{t}$ in Eq. (1) and summand in Eqs. (53) and (54) is $O(d_{1}^{2})$. Since we use these equations from $t=1$ to $t=T+T^{\prime}$, the computational complexity of the RC dynamics is $O(d_{1}^{2}(T+T^{\prime}))=O(d_{1}^{2}T)$. From Eqs. (51) and (52), the computational complexity of the numerator of the IPC is $O(d_{1}^{3})$. Therefore, the computational complexity of a single IPC calculation is $O(d_{1}^{2}T)$. If we evaluate the IPC $N$ times to obtain the sample mean, the computational complexity is $O(Nd_{1}^{2}T)$. As we will empirically see in the next section, $T$ and $T^{\prime}$ should be more than $10d_{1}$. If the data length is taken from $10d_{1}$ to $100d_{1}$ for every $10d_{1}$, the total computational complexity is approximately $10^{3}$ times $O(Nd_{1}^{3})$. Therefore, if $d_{1}$ is increased, it will be difficult to calculate as is. Since the amount of calculation is a trade-off with the estimation accuracy, we first obtain the asymptotic expansion of the variance of the IPC using short data, then determine $N$ and $T$ required for the desired accuracy to prevent excessive calculations. Furthermore, the computation time can be reduced using the bootstrap or jackknife method to decrease the number of samples calculated or by performing parallel computation. However, the computation time cannot be drastically reduced, so $d_{1}\sim 10^{3}$ is expected to be the limit.
Thirdly, our method can be applied to RC systems that satisfy the assumption that the CLT holds for $l(w,U^{\prime}_{t})$, $\nabla l(w,U^{\prime}_{t})$, and $\mu(U^{\prime}_{t})$, the LLN holds for $x(U^{\prime}_{t})x(U^{\prime}_{t})^{\top}$, and $C_{l,\mu}(w)$ is a finite matrix. Thus, for example, if $l(w,U^{\prime}_{t})$, $\nabla_{w}l(w,U^{\prime}_{t})$, or $\mu(U^{\prime}_{t})$ are not ergodic processes, or the mean or covariance matrices are infinite or time-dependent, then Eqs. (41) and (43) cannot be used.